A rectangle is 3 times as long as it is wide and the perimeter is 56. What is the width
Answer by jim_thompson5910(35256)

x = width
3x = length


P = 2L + 2W

56 = 2(3x) + 2x

56 = 6x + 2x

56 = 8x

8x = 56

x = 56/8

x = 7

3x = 3*7 = 21

=======================================================

Width = 7
Length = 21